OpenStudy (speedsterpro):

Taylor complete the problem below. Describe and correct the error in the problem 2^-4*-2*-2*-2*-2=16 A. Taylor changed the negative exponent to a negative base. The problem should be 2^-4=1/2^4=1/2*2*2*2=1/16 B. Taylor didn't multiply correctly. The answer should be 16 C. Taylor made the answer positive. It should be -16 D. Taylor didn't multiply by -1. The problem should be 2^-4=(-1)1/2^4=(-1)1/2*2*2*2=1/16

OpenStudy (mathstudent55):

Didn't I just answer this question?

OpenStudy (mathstudent55):

I think the first * in the problem was supposed to be =.

OpenStudy (mathstudent55):

Problem states: \(2^{-4} = (-2) \times (-2) \times (-2) \times (-2) = 16\) A negative exponent does not mean to make the base negative. A negative exponent changes the power into a fraction according to the rule: \(\Large a^{-n} = \dfrac{1}{a^n} \) Applying the rule of negative exponent, you get \(\Large 2^{-4} = \dfrac{1}{2^4}\) What is \(2^4\) equal to?